OpenStudy (anonymous):

A 16-kg scooter is moving at a speed of 7 m/s. The scooters speed doubles. What is the scooter's kinetic energy when its speed doubles?

rvc (rvc):

\(\rm\huge K.E= \frac{1}{2}\cdot m\cdot v^2\)

rvc (rvc):

if velocity doubles \[\rm K.E=\frac{1}{2}\cdot m\cdot (2v)^2 \\ ~= \frac{4}{2}\cdot m\cdot (v)^2~\\~ =4(\frac{1}{2}\cdot m\cdot (v)^2)\]

OpenStudy (anonymous):

Thnxs

rvc (rvc):

You're welcome so what is the ans?